OpenStudy (lifeisadangerousgame):

Suppose p varies directly as d, and p = 2 when d = 7. What is the value of d when p = 10? Is the answer d = 20/7?

OpenStudy (amistre64):

directly is a p/d = k type relation

OpenStudy (lifeisadangerousgame):

my possible answers are A. d = 35 B. d = 15 C. d = 7/5 D. d = 20/7

OpenStudy (amistre64):

k = 2/7 regardless 10/d = 2/7 id go 70/2 perhaps

OpenStudy (amistre64):

70/2 = 35 :)

OpenStudy (hoblos):

p varies directly as d => p=kd p = 2 when d = 7 => k=2/7 p=2/7 d d= 7/2 p d= 70/2

OpenStudy (lifeisadangerousgame):

Thank you!

OpenStudy (anonymous):

p = 10 * 7/2 = 35